# Entropy-Adaptive Cross-Attention Fusion for Multimodal Speculative Decoding

### v2

@@ Abstract
Speculative decoding (SD) has emerged as a powerful method for accelerating autoregressive generation in large language models (LLMs), yet its integration into vision-language models (VLMs) remains underexplored. We introduce DREAM, a novel speculative decoding framework tailored for VLMs that combines three key innovations: (1) a cross-attention-based mechanism to inject intermediate features from the target model into the draft model for improved alignment, (2) adaptive intermediate feature selection based on attention entropy to guide efficient draft model training, and (3) visual token compression to reduce draft model latency. DREAM enables efficient, accurate, and parallel multimodal decoding with significant throughput improvement. Experiments across a diverse set of recent popular VLMs, including LLaVA, Pixtral, SmolVLM and Gemma3, demonstrate up to 3.6$\times$ speedup over conventional decoding and significantly outperform prior SD baselines in both inference throughput and speculative draft acceptance length across a broad range of multimodal benchmarks. The code is publicly available at: https://github.com/SAI-Lab-NYU/DREAM.git
@@ 1 Introduction
  • Evaluation results show that these three strategies enable DREAM to achieve up to a 3.6×3.6$\times$3.6 × reduction in latency compared to conventional decoding methods across a range of VLMs and tasks, outperforming existing speculative decoding approaches.
@@ 2.1 Speculative Decoding
While the aforementioned techniques focus on SD for language-only models, there has been limited exploration of SD in MLLMs. In speech synthesis, VADUSA li2024fast applies SD to accelerate inference in text-to-speech systems, simultaneously improving synthesis quality. Building on the principles of SD, the authors of raj2024faster propose a multi-token prediction mechanism that significantly boosts inference efficiency for speech generation. In the VLM context, gagrani2024speculative applies speculative decoding to the LLaVA-7B model, demonstrating up to 2.37×2.37$\times$2.37 × speedup by utilizing a lightweight, language-only draft model under memory constraints. IbED leebatch proposes an In-batch Ensemble Drafting strategy that employs ensemble techniques at the batch level without introducing additional model parameters. In contrast, DREAM introduces a novel cross-attention mechanism and adaptive intermediate feature selection to enhance draft model training, resulting in substantial latency reductions.
@@ 2.2 Vision Language Model and Computation Profiling
To quantify the computational cost introduced by visual inputs processing in VLMs, we profile the floating point operations (FLOPs) required by various models, including LLaVA-v1.6-Vicuna-7B liu2023visual , LLaVA-v1.6-Vicuna-7B, Pixtral-12B agrawal2024pixtral , SmolVLM-2B marafioti2025smolvlm , and Gemma3-12B gemmateam2025gemma3technicalreport over the ScienceQA dataset. We select a typical sample which contains a $480\times 300$ image, a prompt with 166 tokens, and use the model to generate 500 tokens, representing an average case across the dataset. FLOPs are measured using PyTorch Profiler. Figure 2 compares the TFLOPs for each model when processing text-only (Txt) versus multimodal (Img+Txt) inputs, revealing an average 2.1×2.1$\times$2.1 × increase in computation with the inclusion of visual data and highlighting the need for more efficient visual processing methods.
@@ 3.2 Adaptive Intermediate Feature Selection for Draft Training
[…]
@@ 4 Empirical Results
We freeze the target VLMs and train only the draft model using the LLaVA mix665k dataset, with 55,000 training samples and 1,000 training samples from each corresponding evaluation benchmark dataset. Training is performed for 68,000 iterations using AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.95$), a learning rate of $3\times 10^{-5}$, and gradient clipping set to 0.5. Each draft model consists of an initial decoder block, an intermediate cross-attention block, and a final decoder block. The weights for the loss terms $\mathcal{L}_{\mathrm{feat}}$, $\mathcal{L}_{\mathrm{intermed}}$, and $\mathcal{L}_{\mathrm{KL}}$ are set to 0.2, 0.2, and 1.0. The parameter sizes are 0.24B for LLaVA-v1.6-Vicuna-7B, 0.37B for LLaVA-v1.6-Vicuna-13B, 0.56B for Pixtral-12B, 0.28B for SmolVLM-2B, and 0.56B for Gemma3-12B. Training is conducted on two NVIDIA A100 80 GB GPUs with batch size set to 4.
During the DREAM evaluation, we use speculative sampling with a batch size of 1, following prior work. The tree structure is configured with $k=4$ child nodes, a depth of 6, and a maximum draft length of 32 tokens. $75\%$ of the visual tokens are retained. These values were chosen based on preliminary experiments to balance aggressive speculation against verification overhead. A full sensitivity analysis and ablation study on these hyperparameters can be found in supplementary materials. All models are evaluated on a single NVIDIA A100 80GB GPU.
@@
LLaVA-v1.6 Vicuna-7B SPD gagrani2024speculative 1.10 1.88 0.81 1.17 1.08 1.87 0.89 1.25 0.91 1.24 1.06 1.76 0.97 1.53
@@ 4.1 Evaluation Results
Table 1 presents the speedup ratios and average acceptance lengths $\tau$ for DREAM compared to baseline methods. Across all tasks and target models, DREAM consistently achieves the highest speedup and longest acceptance lengths. Notably, DREAM delivers a 1.5×1.5$\times$1.5 × to 3.6×3.6$\times$3.6 × speedup over standard autoregressive decoding with the target model, and achieves a $20\%$ to $40\%$ improvement over EAGLE-2. Particularly, we observe that larger models gain greater benefits from DREAM. At $T=0$, DREAM achieves an average speedup of 3.06×3.06$\times$3.06 × on LLaVA-v1.6-Vicuna-7B and 2.65×2.65$\times$2.65 × on Pixtral-12B, whereas the speedups are lower for smaller models, with 2.27×2.27$\times$2.27 × on SmolVLM-2B and 2.23×2.23$\times$2.23 × on LLaVA-v1.6-Vicuna-7B. This is because larger models experience more severe decoding bottlenecks, allowing the draft model to more effectively substitute the costly decoding process of them. Furthermore, models like LLaVA and Pixtral embed visual features directly into the language decoder, offering clearer multimodal cues. This allows DREAM to achieve higher acceptance lengths, for example, $\tau=5.51$ on LLaVA-v1.6-Vicuna-7B. In contrast, models such as Gemma 3-12B, which handle cross-modal information through more complex processing pathways, reach lower acceptance rates, with $\tau=2.36$.
Second, the task-level analysis reveals that long-form QA tasks such as MMT-Bench and ScienceQA benefit most from speculative decoding. These tasks require generating structured answers grounded in high-level visual semantics, which DREAM captures efficiently. For example, on ScienceQA with LLaVA-v1.6-Vicuna-7B, DREAM achieves a speedup of 3.36$\times$ with $\tau=5.32$, outperforming Kangaroo (1.22$\times$) and EAGLE-2 (3.09$\times$). In contrast, OCR-style datasets such as MathVista pose greater challenges due to their reliance on fine-grained character-level recognition, which current draft models struggle to replicate. As a result, all methods show lower $\tau$ and speedup on this benchmark.
Finally, temperature settings critically impact performance. At low temperature ($T=0$), deterministic decoding leads to higher token alignment and longer acceptance spans. For instance, DREAM achieves $\tau=4.52$ on MMT-Bench with Pixtral-12B. When sampling is enabled ($T=1$), performance moderately degrades due to increased token variance, yet DREAM’s tree-based architecture maintains robust performance. On MMT-Bench with Pixtral-12B, DREAM still delivers a speedup of 2.90$\times$ and $\tau=4.02$, outperforming all baselines under the same temperature.
@@ 4.2 Ablation Study
As shown by prior work, tree-structured drafts improve the accepted sequence length compared to chain-structured drafts li2024eagle ; li2024eagle2 ; chen2024sequoia . Figure 4 (b) shows the performance of DREAM with both chain- and tree-structured drafts on the LLaVA-v1.6-Vicuna-7B (LV) and Pixtral-12B (PX) models across the MMT-Bench (MMT) and ScienceQA (SQ) datasets. Tree decoding provides an average of 1.32×1.32$\times$1.32 × speedup on top of our proposed methods described in Section 3. It is important to note that all the baseline algorithms we compared except Kangaroo utilize tree-structured decoding. Even without the help of tree decoding, compared with Kangaroo, DREAM achieves an average of $56\%$ and $75\%$ speedup on LLaVA and Pixtral, respectively.
@@
To improve decoding efficiency, DREAM subsamples and retains only a fraction of visual tokens from the ViT encoder, as detailed in Section 3.3. In Figure 4 (c), we evaluate the effect of different visual token compression ratios, where a fraction of 1 indicates no compression, and 3/4 means $75\%$ of the tokens are kept. Experiments on the LLaVA-v1.6-Vicuna-7B model across MMT-Bench, SEED-Bench, and ScienceQA show that retaining 3/4 of the tokens yields $7\%$ speedup with only a minor reduction in acceptance length. However, when only 1/2 and 1/4 of the tokens are retained, the reduced visual input leads to noticeable information loss, lowering acceptance length and diminishing the speedup. More results are shown in the supplementary results.
Impact of Lambda Settings
5 Conclusion and Limitation
We present DREAM, a speculative decoding framework optimized for VLMs. By integrating visual token compression, cross-attention feature fusion, and adaptive intermediate distillation, DREAM achieves up to 3.6×3.6$\times$3.6 × speedup over standard decoding while maintaining high accuracy, offering a scalable and efficient solution for fast multimodal inference.
Appendix
A1: Tree Based Generation Process
To better illustrate the tree-based generation process, we present a representative example. As shown in Figure 5(a), we begin from a prefix sequence “Fig. Describe :” and a token “It,” which serves as the root node of the draft tree and corresponds to the most recently verified token provided by the target model. Here, the special prefix token “Fig.” denotes the visual tokens representation of the input image. From the root, the draft model expands a tree of candidate tokens across multiple depth layers.
We expand the draft tree by sampling candidate tokens at each depth in parallel. To maintain autoregressive integrity, we apply a tree-based attention mask. Each node in the draft tree only attends to its ancestral path, thereby preventing information leakage across branches.
(a) Tree-based expansion process.
(b) Flattened sequence and its tree-based attention mask.
Figure 5: Tree-based generation and attention mechanism. (a) The draft model generates a tree of candidate tokens starting from a root token (“It”) following the prefix (“Fig. Describe :”). Tokens are expanded layer-by-layer with associated global acceptance scores. Colors denote different roles: purple for prefix tokens, blue for the root, yellow for retained tokens, and gray for pruned branches. (b) The tree-based attention mask ensures that each token attends only to its ancestors in the tree, thereby preserving autoregressive structure and preventing cross-branch leakage.
Top-$k$ Expansion
To prevent exponential expansion, we selectively expand only the top-$k$ tokens with the highest estimated global acceptance scores at each depth layer of the draft tree:
$C_{i}=\prod_{d_{j}\in\text{Path(root},d_{i})}c_{j}$
[…]
Reranking Strategy
To mitigate the verification cost associated with a large draft tree, we use the reranking step that prunes low-confidence branches. Since tokens at deeper levels tend to have lower acceptance probabilities, we rerank all candidate tokens across all depths based on their global scores and select the top-$m$ tokens for verification.
To preserve structural consistency, we retain all ancestors of selected tokens and resolve ties in favor of shallower nodes. The selected subtree is then flattened into a linear sequence for verification. During this phase, we continue to apply the tree-based attention mask, ensuring that each token can only attend to its ancestral tokens, thus preserving autoregressive constraints. In Figure 5(a), red-bordered tokens represent those selected during the reranking process with top-$m=6$, including the root node. Figure 5(b) visualizes the flattened sequence and its corresponding attention mask.
A2: Implementation Details
All models are based on their official Hugging Face implementations. Baseline methods are executed using their publicly released code and default configurations, with minimal modifications to support VLM inputs. During inference, all models are run with KV caching enabled to ensure efficient autoregressive decoding.
For draft model training of the baseline models, including Kangaroo liu2024kangaroo , Medusa cai2024medusa , EAGLE li2024eagle and EAGLE-2 li2024eagle2 , we follow their publicly available training scripts, introducing only minimal adjustments to accommodate VLM input formats. To ensure a fair comparison, all methods are trained and evaluated using the same dataset, number of training epochs, learning rate, batch size, and hardware environment. We fix random seeds across all runs to reduce performance variance and report the average over three runs.
Table 3: Ablation Study: Effect of Visual Token Compression Rate on Speedup and Token Acceptance
@@ 5 Conclusion and Limitation
A3: VTC Sensitivity Analysis
We analyze the impact of the visual token compression (VTC) rate on both inference speed and draft token accuracy within our pipeline. We experiment with four compression rates of $100\%$, $75\%$, $50\%$, and $25\%$, and evaluate their effects using two metrics: speedup ratio ($S$) and average accepted token length ($\tau$), across three benchmarks: MMT-Bench, SEED-Bench, and ScienceQA.
As shown in Table 3, the data show that a 75% VTC rate is usually the sweet spot. For LLaVA-13B and Pixtral-12B it delivers the best or second-best speedup (3.68$\times$ and 2.93$\times$ on MMT-Bench) while keeping the average accepted length $\tau$ within  2% of the full-token baseline. Cutting the visual tokens further to 50% or 25% yields diminishing speed gains and drives $\tau$ down by 10–25%, indicating that verification rejects more draft tokens once too much visual context is removed.
A4: Statistical Significance Analysis
We evaluate the statistical significance of DREAM’s improvements over six baseline methods on two key metrics—speedup ratio ($S$) and average accepted token length ($\tau$)—across six datasets (MMT-Bench ying2024mmtbench , SEED-Bench-2 li2023seed2 , ScienceQA lu2022learn , OCRBench Liu_2024 , ChartQA masry-etal-2022-chartqa , and MathVista lu2024mathvista ), using LLaVA-v1.6-Vicuna-13B liu2023improved at temperature $T=0$.
DREAM consistently achieves statistically significant improvements in speedup ratio ($S$) across all baselines, with large margins and strong effect sizes. For average accepted token length ($\tau$), DREAM also shows significant improvements in nearly all comparisons, demonstrating its ability to not only accelerate inference but also maintain high output quality.
These results highlight DREAM’s robustness across diverse datasets and baselines, offering both faster decoding and better verification efficiency under strong backbone models.
Table 4: Significance test of DREAM vs. baselines, with Bonferroni-corrected $t$-tests ($\alpha=0.0083$)
Method $\Delta S$ $\Delta\tau$
Mean 95% CI $p$ Mean 95% CI $p$
SPD 2.09 [1.66, 2.52] 0.0001 3.36 [3.01, 3.72] 0.0000
Kangaroo 1.70 [1.17, 2.23] 0.0004 3.15 [2.46, 3.84] 0.0001
Medusa 1.32 [1.02, 1.63] 0.0001 2.10 [1.41, 2.78] 0.0005
Hydra 1.19 [0.91, 1.47] 0.0001 1.88 [1.23, 2.53] 0.0007
EAGLE 0.96 [0.68, 1.25] 0.0003 1.46 [0.95, 1.97] 0.0007
EAGLE-2 0.79 [0.56, 1.02] 0.0005 0.42 [0.05, 0.79] 0.0280

### v3

@@ Abstract
[…]
@@ 1 Introduction
  • Evaluation results show that these three strategies enable DREAM to achieve up to a $3.6\times$ reduction in latency compared to conventional decoding methods across a range of VLMs and tasks, outperforming existing speculative decoding approaches.
@@ 2.1 Speculative Decoding
While the aforementioned techniques focus on SD for language-only models, there has been limited exploration of SD in MLLMs. In speech synthesis, VADUSA li2024fast applies SD to accelerate inference in text-to-speech systems, simultaneously improving synthesis quality. Building on the principles of SD, the authors of raj2024faster propose a multi-token prediction mechanism that significantly boosts inference efficiency for speech generation. In the VLM context, gagrani2024speculative applies speculative decoding to the LLaVA-7B model, demonstrating up to $2.37\times$ speedup by utilizing a lightweight, language-only draft model under memory constraints. IbED leebatch proposes an In-batch Ensemble Drafting strategy that employs ensemble techniques at the batch level without introducing additional model parameters. In contrast, DREAM introduces a novel cross-attention mechanism and adaptive intermediate feature selection to enhance draft model training, resulting in substantial latency reductions.
@@ 2.2 Vision Language Model and Computation Profiling
To quantify the computational cost introduced by visual inputs processing in VLMs, we profile the floating point operations (FLOPs) required by various models, including LLaVA-v1.6-Vicuna-7B liu2023visual , LLaVA-v1.6-Vicuna-7B, Pixtral-12B agrawal2024pixtral , SmolVLM-2B marafioti2025smolvlm , and Gemma3-12B gemmateam2025gemma3technicalreport over the ScienceQA dataset. We select a typical sample which contains a $480\times 300$ image, a prompt with 166 tokens, and use the model to generate 500 tokens, representing an average case across the dataset. FLOPs are measured using PyTorch Profiler. Figure 2 compares the TFLOPs for each model when processing text-only (Txt) versus multimodal (Img+Txt) inputs, revealing an average $2.1\times$ increase in computation with the inclusion of visual data and highlighting the need for more efficient visual processing methods.
@@ 3.2 Adaptive Intermediate Feature Selection for Draft Training
where $\mathrm{smoothL1}(x,y)$ equals $\dfrac{1}{2}(x-y)^{2}$ if |xy|<1|x-y|<1, and $|x-y|-\dfrac{1}{2}$ otherwise girshick2015fast .
@@ 4 Empirical Results
We freeze the target VLMs and train only the draft model using the LLaVA mix665k dataset, with 55,000 training samples and 1,000 training samples from each corresponding evaluation benchmark dataset. For benchmarks without predefined train–test splits (MMT-Bench, SEED-Bench-2, MathVista, and OCRBench), we applied stratified random sampling using train_test_split (with random_state=42 for reproducibility) to select 1,000 training samples and up to 3,000 test samples per dataset. For ChartQA and ScienceQA, which already provide official splits, we randomly sampled 1,000 training examples from their respective training sets. All selected training samples were then processed through the target model to generate ground-truth responses in conversational form.
Training is performed for 68,000 iterations using AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.95$), a learning rate of $3\times 10^{-5}$, and gradient clipping set to 0.5. Each draft model consists of an initial decoder block, an intermediate cross-attention block, and a final decoder block. The weights for the loss terms $\mathcal{L}_{\mathrm{feat}}$, $\mathcal{L}_{\mathrm{intermed}}$, and $\mathcal{L}_{\mathrm{KL}}$ are set to 0.2, 0.2, and 1.0. The parameter sizes are 0.65B for LLaVA-v1.6-Vicuna-7B, 0.98B for LLaVA-v1.6-Vicuna-13B, 0.9B for Pixtral-12B, 0.28B for SmolVLM-2B, and 0.9B for Gemma3-12B. Training is conducted on two NVIDIA A100 80 GB GPUs with batch size set to 4. Our training procedure adopts a two-stage pipeline: (1) offline calibration for feature selection, and (2) draft model training. During the calibration stage, we compute the attention entropy for each sample once to identify the optimal intermediate layer $\ell^{*}$, and cache these selections for subsequent training.
During the DREAM evaluation, we use speculative sampling with a batch size of 1, following prior work. The tree structure is configured with $k=4$ child nodes, a depth of 6, and a maximum draft length of 32 tokens. Note that the average accepted token length ($\tau$) reported in Table 1 may exceed a depth of 6, as the target model generates an additional token once all drafts are verified, resulting in up to 7 tokens per round leviathan2023fast . $75\%$ of the visual tokens are retained. These values were chosen based on preliminary experiments to balance aggressive speculation against verification overhead. A full sensitivity analysis and ablation study on these hyperparameters can be found in supplementary materials. All models are evaluated on a single NVIDIA A100 80GB GPU.
All models are based on their official Hugging Face implementations. Baseline methods, including Kangaroo liu2024kangaroo , Medusa cai2024medusa , EAGLE li2024eagle and EAGLE-2 li2024eagle2 , are executed using their publicly released code and default configurations, with minimal modifications to support VLM inputs. During inference, all models are run with KV caching enabled to ensure efficient autoregressive decoding. To ensure a fair comparison, all methods are trained and evaluated using the same dataset, number of training epochs, learning rate, batch size, and hardware environment. We fix random seeds across all runs to reduce performance variance and report the average over three runs.
@@
LLaVA-v1.6 Vicuna-7B
SPD gagrani2024speculative 1.10 1.88 0.81 1.17 1.08 1.87 0.89 1.25 0.91 1.24 1.06 1.76 0.97 1.53
@@ 4.1 Evaluation Results
Table 1 presents the speedup ratios and average acceptance lengths $\tau$ for DREAM compared to baseline methods. Across all tasks and target models, DREAM consistently achieves the highest speedup and longest acceptance lengths. Notably, DREAM delivers a $1.5\times$ to $3.6\times$ speedup over standard autoregressive decoding with the target model, and achieves a $20\%$ to $40\%$ improvement over EAGLE-2. Particularly, we observe that larger models gain greater benefits from DREAM. At $T=0$, DREAM achieves an average speedup of $3.06\times$ on LLaVA-v1.6-Vicuna-7B and $2.65\times$ on Pixtral-12B, whereas the speedups are lower for smaller models, with $2.27\times$ on SmolVLM-2B and $2.23\times$ on LLaVA-v1.6-Vicuna-7B. This is because larger models experience more severe decoding bottlenecks, allowing the draft model to more effectively substitute the costly decoding process of them. Furthermore, models like LLaVA and Pixtral embed visual features directly into the language decoder, offering clearer multimodal cues. This allows DREAM to achieve higher acceptance lengths, for example, $\tau=5.51$ on LLaVA-v1.6-Vicuna-7B. In contrast, models such as Gemma 3-12B, which handle cross-modal information through more complex processing pathways, reach lower acceptance rates, with $\tau=2.36$.
[…]
@@ 4.2 Ablation Study
As shown by prior work, tree-structured drafts improve the accepted sequence length compared to chain-structured drafts li2024eagle ; li2024eagle2 ; chen2024sequoia . Figure 4 (b) shows the performance of DREAM with both chain- and tree-structured drafts on the LLaVA-v1.6-Vicuna-7B (LV) and Pixtral-12B (PX) models across the MMT-Bench (MMT) and ScienceQA (SQ) datasets. Tree decoding provides an average of $1.32\times$ speedup on top of our proposed methods described in Section 3. It is important to note that all the baseline algorithms we compared except Kangaroo utilize tree-structured decoding. Even without the help of tree decoding, compared with Kangaroo, DREAM achieves an average of $56\%$ and $75\%$ speedup on LLaVA and Pixtral, respectively.
@@
To improve decoding efficiency, DREAM subsamples and retains only a fraction of visual tokens from the ViT encoder, as detailed in Section 3.3. In Figure 4 (c), we evaluate the effect of different visual token compression ratios, where a fraction of 1 indicates no compression, and 3/4 means $75\%$ of the tokens are kept. Experiments on the LLaVA-v1.6-Vicuna-7B model across MMT-Bench, SEED-Bench, and ScienceQA show that retaining 3/4 of the tokens yields $7\%$ speedup with only a minor reduction in acceptance length. However, when only 1/2 and 1/4 of the tokens are retained, the reduced visual input leads to noticeable information loss, lowering acceptance length and diminishing the speedup.
Table 3: Effect of Visual Token Compression Rate on Speedup and Token Acceptance
@@
Specifically, we examine how the visual token compression (VTC) rate affects both inference speed and draft token accuracy in DREAM. We test four compression levels (100%, 75%, 50%, and 25%) and evaluate their effects using two key metrics: the speedup ratio ($S$) and the average accepted token length ($\tau$) across three benchmarks: MMT-Bench, SEED-Bench, and ScienceQA.
[…]
Impact of Lambda Settings
5 Conclusion and Limitation
We present DREAM, a speculative decoding framework optimized for VLMs. By integrating visual token compression, cross-attention feature fusion, and adaptive intermediate distillation, DREAM achieves up to $3.6\times$ speedup over standard decoding while maintaining high accuracy, offering a scalable and efficient solution for fast multimodal inference.